OpenStudy (katherinesmith):

Find the value of cos210°

OpenStudy (katherinesmith):

@KeithAfasCalcLover

OpenStudy (anonymous):

We know that: \[cos(x)=-cos(x-180^\circ)\] So: \[cos(210^\circ)=-cos(210^\circ-180^\circ)=-cos(30^\circ)\] We need the special triangle:  So therefore: \[cos(210^\circ)=-cos(30^\circ)=-\left(\frac{\sqrt{3}}{2}\right)=-\sqrt{1.5}\]